Z^3=pqxy by jacobi's method


1
Expert's answer
2021-01-26T04:21:31-0500

p=-u1/u3; q=-u2/u3

z3=u1u2xy/u32

z3u32 - u1u2xy=0

"\frac{dx}{-u_2xy}=\frac{dy}{-u_1xy}=\frac{dz}{2z^3u_3}=\frac{du_1}{u_1u_2y}=\frac{du_2}{u_1u_2x}=\frac{du_3}{-3z^2u_3^2}"

u1dx=u2dy=a

"\frac{dz}{2z}= \frac{du_3}{-3u_3}"

"\frac{log(z)}{2}= \frac{log(u_3)}{-3}"

u3=ze-3/2+b

u=2a+(ze-3/2 + b)dz=2a+z2e-3/2/2+bz+c

u=c

2a+z2e-3/2/2+bz=0

"z=\frac{-b^+_- \sqrt{b^2-4ae^{-3/2}}}{e^{-3/2}}"